40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 81 of 327  Not logged in ELOG logo
ID Date Author Type Category Subject
  12434   Tue Aug 23 19:35:38 2016 gautamUpdateSUSETMY back in IFO

[johannes, gautam]

Summary: Today we moved the suspended ETMY optic back into the chamber from the cleanroom. Once in the chamber, we positioned the optic using the stops that marked the previous position of the optic. We then shortened the arm length by 19mm (in order to match the X and Y arm lengths. The F.C. coat on the HR face was removed prior to the final placement of the optic. We then adjusted the OSEM positions in their holders to get the sensor outputs to half their maximum value.

We did not get to check where the input beam hits the optic or see if the pitch balance of the optic is such that the reflected beam makes it back to the ITM. The plan for tomorrow is to do this. 


Part 1: Cleanroom work

  • We worked a little more on trying to adjust the rotational position of the OSEM coils in order to minimize the coupling of the bounce mode into the sensor signals. 
  • We had limited success in this regard. After about an hour, we concluded that it made more sense to do this in the chamber itself. For one thing, the drive electronics for the Y end are different (in the cleanroom, we are using the X end electronics, satellite box etc.).
  • We adjusted the position of the OSEMs till the sensor output readout was half the open value as best as we could. We also made sure that the wire was in the groove on both sides and that the magnets were well centered in the vertical direction relative to the OSEM coils and that there was no danger of knocking any magnets off (see attached pictures).
  • We then engaged all the EQ stops, and transferred the suspension cage to a cart (topped with Al foil, wiped clean) for transportation to the Y-end (with OSEMs left in).

Part 2: Transportation of optic

  • Nothing special here, just took great care while going over bumps near doors between the cleanroom and the IFO, and along the Y-arm itself.
  • Definitely a 2 man job - one person can lift a pair of wheels over any bumps while the other can make sure there is no danger of the cage toppling over. 

Part 3: Chamber work

  • PSL shutter was closed for this part of work. Earlier today, I found that C1SUSAUX had failed yet again (why are all the slow computers dying more often nowadays?!). I restarted the slow machine, and locked the mode cleaner. The alignment hadn't drifted so much from when EricQ had last aligned the IMC, and with only minimal tweaking, I was able to lock the IMC and see a beam on the REFL camera.
  • First, I transferred the suspension cage onto the edge of the table inside the chamber. Care was taken not to accidentally place the cage onto the trailing OSEM wires.
  • There were some specks of dust on the barrel of the optic, and also the cage. These were removed with clean wipes and isopropanol.
  • I judged that it would be too precarious to remove the F.C. with the optic in its final desired position. So we decided to take the coat off with the optic at the edge of the table. The central part of the HR face looks pretty clean. Even though the whole HR face was cleaned with F.C., the part that was left uncovered prior to putting the optic back into the chamber has a few specks of dust on it (see attachments). These could not be removed just by blowing ionized air. I was hesitant to drag wipe the optic, so I left things as is. In any case, the optic as a whole is MUCH cleaner (to my eye at least) than prior to the cleaning. 
  • Conveniently, the stops marking the previous position of the optic were on the far side and back.
  • Since we wanted to shorten the Y arm length by 2 cm, we placed a clean steel ruler of width 19mm in front of the rear stop (see attached pictures). I then moved the cage back along the side stop till I hit the ruler.
  • I then clamped the optic down, removed the spacing ruler, and re-adjusted the position of the rear stop to mark the new position of ETMY.
  • We were concerned that the change of position of the cage on the table affected the leveling. Checking with a clean spirit level, we found evidence of a slight tilt in the direction towards the vertex of the IFO, as expected from the way the ETMY cage was moved. To compensate for this, I moved one of the counterweight masses (see attachments) till the spirit level showed the table to be level (to its resolution) in two perpendicular directions
  • We then plugged in the OSEMs into the DB25 connectors on the table. We found that the Y-end electronics were giving different readouts from what we had been seeing in the cleanroom with the X end electronics (not surprising I guess). We resolved to pull out all the OSEMs, check their maximum sensor output values, and re-insert them till the sensor output was half this maximum as best as we could. NOTE TO SELF: UPDATE THE WIKI PAGE!
  • We turned on the damping, and found that the exisiting input matrix performs fairly well.
  • We took a quick look at the spectra of the sensor outputs - interestingly, with the suspension on the seismic stacks inside the chamber, the 16.4 Hz bounce mode peak showed up clearly (these were totally absent in the cleanroom). I did not attempt any fine rotation of the OSEMs in the holders (it is not even clear to me how good/bad the present configuration is) because I reasoned we first need to apply a pitch bias to get the beam back to the ITMY chamber and then re-adjust the OSEM coils. The bounce mode decoupling will be the last step. 
  • For tonight, we decided to leave the optic freely swinging (with EQ stops close by) so that tomorrow, we can look at the offline spectra of sensor outputs and if necessary, re-diagonalize the suspension. 
  • After checking nothing unwanted was left behind in the chamber, we closed it up for tonight.

Plan for tomorrow:

  • Pitch balancing check (by looking at reflected beam at ITMY)
  • Re-adjust OSEMs on ETMY, minimize bounce mode coupling into sensor outputs
  • Make Y arm cavity by re-positioning ITMY

Attachments:

Attachment #1: Wire is in groove in side without OSEM

Attachment #2: Wire is in groove in side with OSEM (picture taken with OSEM coil removed)

Attachment #3: UL magent relative to OSEM coil

Attachment #4: LL magent relative to OSEM coil

Attachment #5: LR magnet relative to OSEM coil

Attachment #6: UR magnet relative to OSEM coil

Attachment #7: Side magnet relative to OSEM coil

Attachment #8: ETMY HR face with F.C. film removed. Non-covered part isn't super clean, but the covered part itself does not have any large specks of dust visible.

Attachment #9: Scheme adopted to shorten Y arm length by 19mm.

Attachment #10: Current situation inside EY chamber. Counterweight that was moved to balance the table is indicated.

 

Attachment 1: IMG_3025.JPG
IMG_3025.JPG
Attachment 2: IMG_3035.JPG
IMG_3035.JPG
Attachment 3: IMG_3030.JPG
IMG_3030.JPG
Attachment 4: IMG_3029.JPG
IMG_3029.JPG
Attachment 5: IMG_3028.JPG
IMG_3028.JPG
Attachment 6: IMG_3027.JPG
IMG_3027.JPG
Attachment 7: IMG_3026.JPG
IMG_3026.JPG
Attachment 8: IMG_3036.JPG
IMG_3036.JPG
Attachment 9: IMG_3038.JPG
IMG_3038.JPG
Attachment 10: IMG_3045.JPG
IMG_3045.JPG
  12433   Tue Aug 23 17:05:20 2016 PrafulUpdateElectronicsSoldered Circuit Working

I remade another soldered circuit, adding extra 100uF electrolytic bypass capacitors at the input and output of the voltage regulator and ensuring that every grounded component now has its own path to ground rather than going through other elements. This circuit now seems to be working just like the solderless circuit. Attached is the transfer function of the soldered circuit, which matches with the result from the solderless circuit.

 

soldered_transfer_function.png

solderless_transfer_function.png

Here are both on the same figure- they are about overlapping but are slightly different if you zoom in enough.

both_transfer.png

I have also attached a new version of the circuit schematic to reflect the changes and to make the physical layout more clear.

simple_ampv2.pdf

My next step for these last few days this summer will be designing a PCB using Altium. I've emailed Varun about how to use Altium on the iMac but he hasn't responded. If anyone else knows how to use the software, please let me know.

Attachment 2: soldered_transfer_function.png
soldered_transfer_function.png
Attachment 3: soldered_transfer_function.png
soldered_transfer_function.png
Attachment 5: solderless_transfer_function.png
solderless_transfer_function.png
Attachment 6: both_transfer.png
both_transfer.png
Attachment 8: both_transfer.png
both_transfer.png
Attachment 10: simple_ampv2.pdf
simple_ampv2.pdf
  12432   Tue Aug 23 09:50:17 2016 Max IsiUpdateGeneralSummary pages down due to cluster maintenance

Summary pages down today due to schedulted LDAS cluster maintenance. The pages will be back automatically once the servers are back (by tomorrow).

  12431   Mon Aug 22 18:35:16 2016 PrafulUpdatePEMthe lab temp is up

The temperature is decreasing slowly but is still above 24 C.

temp_plot.png

Quote:

The IFO room temp is up a bit and it is coming down. The out side temp is not really high.

 

Attachment 1: temp_plot.png
temp_plot.png
Attachment 3: temp_plot.png
temp_plot.png
  12430   Mon Aug 22 18:04:24 2016 gautamUpdateSUSETMY OSEMs inserted

[Johannes, gautam]

We worked on trying to insert the OSEMs in the optimal positions such that the coupling of the bounce mode into the OSEM sensor signals was minimised.

First, I gave the barrel of the optic a wipe with some optical tissue + acetone in order to remove what looked like some thin fibres of dried first contact. It may be that while I was applying the F.C., the HEPA air flow deposited these on the barrel. In any case, they came off easily enough. There is still a few specks of dust on various parts of the barrel, but it is likely that these can just be removed with the ionized air jet, which we can do after putting the optic in the chamber.

We then did the usual OSEM insertion till the magnets neutral position was such that the sensor output was ~50% of the fully open value (turned the HEPA off for the remainder of this work). I tweaked the bottom OSEM plate a little in order to center the magnets relative to the coil as best as possible. Once this was done, we attempted to look at spectra of the sensor outputs, with 0.05 Hz bandwidth - however, we were unable to identify any peak at 16.4 Hz, which is what a Jan 2015 measured value wiki page claims the bounce mode frequency is (although this was an in vacuum measurement). There were a couple of peaks at ~15.7 Hz and ~16.7 Hz, but I can't think of any reason why the bounce mode resonance should have changed so much - after all, this is ETMY for which no standoff regluing was done. The only difference is that there is some first contact + peek mesh on the HR face now, but I doubt this can modify the bounce resonance frequency so much (this is just my guess, I will have to back this up with a calculation).

Anyways we decided to take this up again tomorrow. Things are progressing fairly well now, I hope to be able to put in ETMY back into the chamber at some point tomorrow and commence re-alignment of the interferometer. I've left the OSEMs in for today, with the EQ stops not engaged but close by. HEPA has been turned back on.

  12429   Mon Aug 22 16:33:32 2016 SteveUpdatePEMthe lab temp is up

The IFO room temp is up a bit and it is coming down. The out side temp is not really high.

Attachment 1: temp.png
temp.png
  12428   Mon Aug 22 13:06:11 2016 gautamUpdateSUSETMY suspended

Today morning, I suspended ETMY and made the same checks dscribed below. The clamping went smoothly, 5 in. lb. of torque seems sufficient, in the limited observation time, there has been no evidence of wire sag. Today afternoon, we will go about putting the OSEM coils in, setting their equilibrium points etc. This may need to be re-done once the optic is in the chamber and the first contact has come off, but at least we can coarsely place them in the relative convenience of the cleanroom. 

GV EDIT 9.15pm 22 Aug: Eric had a look at both towers and pointed out that I had neglected to use washers on the wire stops. After consultation with Steve, I decided that it is not worth it to remove the clamp and re-suspend the optic - it is likely that the current suspension process will have caused new grooves in the suspension block, which will have to be removed, and the sanding process did not work so well last time. In any case, the net effect of this will be that the actual torque with which the clamp is tightened will be slightly different from 5 in. lb., but since there is no evidence that the clamp isn't tight enough / is too tight, I think it is okay to push ahead. 

Quote:
  • ETMX has been successfully suspended
  • I've used one of the new wire clamps, and also the new suspension wire 
  • Because the HR face has first contact, pitch balancing cannot be checked at this point. But since the pitch balance was checked after the standoff was glued, there is no reason to believe it would have changed
  • Heights of the two scribe lines were checked with the microscope and verified to be at 5.5" above the tabletop. Also checked the position of the scribe line on the bottom of the optic to make sure the optic wasn't somehow rotated
  • Checked that wire was in the groove in the standoff on both sides, and that the optic was freely hanging with no EQ stops engaged. I also verified that there are no obvious kinks/other funny features where the wire is in contact with the optic barrel below the standoffs.
  • Wire clamps were tightened with the new torque wrench and 5 in. lb. (0.56 N m) of torque. Primary clamp was successfully tightened. However, the wire snapped between the primary and secondary clamps on one side. It is unclear to me how or why this happened. But since the primary wire clamp is the important one, I don't think it is worth re-suspending ETMX all over again
  • I've left the cage on the flow bench for now, with EQ stops engaged. OSEM coils have yet to be inserted, but I suppose we want to do this in the vacuum chamber now to do the fine rotation to minimize the bounce mode in the OSEM signals
  • I've prepared ETMY and its cage for suspension, will work on it tomorrow

 

Attachment 1: IMG_3019.JPG
IMG_3019.JPG
  12427   Sun Aug 21 17:21:22 2016 PrafulUpdateElectronicsProblems with PCB Circuit

For the past week, I've been trying to make a soldered amplifier circuit to use in a prototype box, However, I've been running into this same issue. The circuit, pictured below, works fine on a solderless breadboard.

simple_amp.png

When I amplify a sine wave, I get a clean looking result at the output on the solderless breadboard:

However, on my soldered circuit, if I turn up the negative voltage supply from the power supply past about -12.5V (the target is -15V), I get a strange signal that Gautam suggested looks like some kind of discharging.

At -12.3 V (soldered breadboard):

At -15.0 V (soldered breadboard):

The signal is much noisier. Zooming in on this second signal, this pattern appears:

This pattern is also showing up even when there is no input from the function generator and the circuit is just given a voltage supply of +/- 15V:

I have tried switching out both the positive and negative voltage regulators, the opamp, and remaking and resoldering the entire circuit but I'm still getting the same signal, which is absent from the solderless circuit. This output was produced with a function generator, so I have also ruled out the microphone as a source of this extra noise. The voltage dependence of this problem made me think it was the voltage regulator, but I've switched out the voltage regulator multiple times and it's still showing up. I'm not sure why this signal appears only as the negative voltage supply is increased- there is no problem with increasing the positive input voltage. Please let me know if you have any ideas as to what component or issue could be causing this.

Attachment 2: simple_amp.png
simple_amp.png
Attachment 4: clean.jpg
clean.jpg
Attachment 5: -12.jpg
-12.jpg
Attachment 6: -15.jpg
-15.jpg
Attachment 7: pat1.jpg
pat1.jpg
Attachment 8: pat2.jpg
pat2.jpg
Attachment 10: bad.jpg
bad.jpg
Attachment 11: pattern.jpg
pattern.jpg
Attachment 12: pattern2.jpg
pattern2.jpg
Attachment 13: pat2.jpg
pat2.jpg
Attachment 16: patternzoomed.jpg
patternzoomed.jpg
  12426   Sun Aug 21 16:23:05 2016 gautamUpdateSUSETMX suspended
  • ETMX has been successfully suspended
  • I've used one of the new wire clamps, and also the new suspension wire 
  • Because the HR face has first contact, pitch balancing cannot be checked at this point. But since the pitch balance was checked after the standoff was glued, there is no reason to believe it would have changed
  • Heights of the two scribe lines were checked with the microscope and verified to be at 5.5" above the tabletop. Also checked the position of the scribe line on the bottom of the optic to make sure the optic wasn't somehow rotated
  • Checked that wire was in the groove in the standoff on both sides, and that the optic was freely hanging with no EQ stops engaged. I also verified that there are no obvious kinks/other funny features where the wire is in contact with the optic barrel below the standoffs.
  • Wire clamps were tightened with the new torque wrench and 5 in. lb. (0.56 N m) of torque. Primary clamp was successfully tightened. However, the wire snapped between the primary and secondary clamps on one side. It is unclear to me how or why this happened. But since the primary wire clamp is the important one, I don't think it is worth re-suspending ETMX all over again
  • I've left the cage on the flow bench for now, with EQ stops engaged. OSEM coils have yet to be inserted, but I suppose we want to do this in the vacuum chamber now to do the fine rotation to minimize the bounce mode in the OSEM signals
  • I've prepared ETMY and its cage for suspension, will work on it tomorrow
Attachment 1: IMG_3018.JPG
IMG_3018.JPG
  12424   Fri Aug 19 22:51:12 2016 gautamUpdateSUSETMs first-contacted

I've applied first contact to both the ETMs. They're now ready to be suspended. I've also cut up some lengths of the new wire and put them in the oven for a 12 hour 70C bake. 

  • For both ETMs, I first applied first contact to the bulk of the HR and AR surfaces (all the way out to the edge for the HR, for the AR as large an area as possible without getting too close to the magnets). Calum recommended pouring first contact onto the horizontal optic, but since I had no practise with this method, I opted not to try it out for the first time on our ETMs
  • After allowing this to dry for 24 hours, I peeled this layer off. Visual inspection suggests that the whole film came off cleanly. 
  • I then applied first contact to a smaller area around the center of the optic for only the HR surface. This will only be peeled off once the suspended optic is back in the vacuum chamber. This way, we keep the HR face protected for as long as possible.
  • Even though we applied F.C to both faces of the ITMs, I don't think its so important to keep a film on the AR side of the ETMs till we take it in. So I didnt re-coat the AR side with a smaller area of F.C. This way, if we want, we can do the OSEM assembly in the cleanroom without having to worry about peeling the F.C off with limited access to the rear of the optic.
  • I also opted to bake some lengths of the newly arrived steel wire for suspension. Not sure how important/useful this bake will be.

Unless we want the AR surface to also have a small F.C coat until the optic is in the vacuum chamber, I think I will proceed with re-suspending the ETMs..

  12423   Thu Aug 18 15:16:09 2016 SteveUpdateSUSSOS sus wire is in

 

Stress Relieved 0.0017"  Music Wire  CFW P/N: CFW2035025,   Made 08-17-2016                                

    Old  2003    New  2016  
GBL 358.9 240.610 grams
UTS 357,061 229,603 PSI
YTS 343,211 177,371 PSI
ELONG 2.38 0.8 %
HEAT 10622 10622  

GBL (grams breaking load )         

UTS (ultimate tensile strength)                                 

YTS (yield tensile strength) 

ELONG (elongation)  

Quote:

0.0017" OD., 500ft steel music wire ordered. Pictures of the existing roll are below. It will be on 8" OD. spool too.

 

 

Attachment 1: 0.0017new.jpg
0.0017new.jpg
  12422   Thu Aug 18 14:14:20 2016 gautamUpdateSUSAir-bake of towers - finished

I took the two cages, wires and wire clamps out this morning, back into the cleanroom after their 12 hour 70C bake. 

I've also applied first contact to the AR face of the optics. Steve is preparing a jig which will allow us to apply first contact on the HR side with the optic horizontal. The idea is to apply a large coating first, to clean the bulk of the HR surface, and peel it off before re-suspending the optic. Then we can paint on a smaller area, suspend the optic (and hope the pitch balancing is alright) before taking the whole assembly into the chamber where it will be peeled off. 

Calum recommended that we buy a new ionizing gun + electrometer assembly (apparently our current set up is woefully obsolete) but I don't know if we can have these in time for the first contact peeling...

  12421   Thu Aug 18 08:17:16 2016 SteveUpdateSUSwire clamping preparation

The wire inprints were removed by 800P grain paper [Norton 73568] The SS bridge block now has an undesireble vally in the wire location.

The sus bridges were soaked in acetone over night and sonicated to remove residual sand paper.

 

Quote:

I just put in the following into the air bake oven for a 12 hour, 70C bake:

  • ETMX and ETMY cages (with sanded suspension blocks loosely tightened for now, we will tighten them after the bake)
  • 13 new wire clamps that were recently made by the shop
  • 7 lengths of suspension wire (since the new wire is unlikely to arrive for another 2 weeks). This should be sufficient in case we overtighten the wire clamps a couple of times and the wire snaps.

I put these in at 10.30pm. So the oven will be turned off at 10.30am tomorrow morning. The oven temperature seems stable in the region 70-80 C (there is no temperature control except for the in built oven control, I just adjusted the dial till I found the oven remains at ~70C.

Tomorrow, we will look to put on first contact onto the ETMs, and then get about to re-suspending them.

 

Attachment 1: 800P.jpg
800P.jpg
  12420   Wed Aug 17 23:00:57 2016 gautamUpdateSUSAir-bake of towers

I just put in the following into the air bake oven for a 12 hour, 70C bake:

  • ETMX and ETMY cages (with sanded suspension blocks loosely tightened for now, we will tighten them after the bake)
  • 13 new wire clamps that were recently made by the shop
  • 7 lengths of suspension wire (since the new wire is unlikely to arrive for another 2 weeks). This should be sufficient in case we overtighten the wire clamps a couple of times and the wire snaps.

I put these in at 10.30pm. So the oven will be turned off at 10.30am tomorrow morning. The oven temperature seems stable in the region 70-80 C (there is no temperature control except for the in built oven control, I just adjusted the dial till I found the oven remains at ~70C.

Tomorrow, we will look to put on first contact onto the ETMs, and then get about to re-suspending them.

Attachment 1: IMG_3006.JPG
IMG_3006.JPG
  12419   Wed Aug 17 22:09:04 2016 ranaUpdateSUSSOS sus wire ordered

Not really true that it passed. That's just an arbitrary margin. Best to throw away all the old wire. We have no quantitative estimate of what the real torque should be. Its just feelings.

Quote:

The wire will arrive in 1-2 weeks. It is a new production. Brad Snook of Ca Fine Wire was suprised that we are still using the 13 years old wire.  Oxidation is an issue with iron contained steel wire.

He would not give me a shelf life time on it. He recommended to check the strenght of it before usage. It passed with safety factor of 2 just recently.

In the future we'll store the new spool in oxigen free nitrogen environment..

 

  12418   Wed Aug 17 16:28:46 2016 KojiUpdateCOCRC folding mirrors - Numerical review

For the given range of the PR3/SR3 RoCs for both cases, all the resulting numbers such as TMSs/mode matching ratios look reasonable to me.

  12417   Wed Aug 17 14:37:36 2016 gautamUpdateCOCRC folding mirrors - Numerical review
Quote:

 

Cavity One-way Gouy phase [rad]           TMS [MHz]           
PRX 0.244 1.730
PRY 0.243 1.716
SRX 0.197 1.743
SRY 0.194 1.717

So, there are regions in parameter space for both options (i.e. keep current G&H mirrors, or order two new sets of folding mirrors) that get us close to the design numbers...

Keeping these design numbers in mind, here are a few possible scenarios. The "designed" TMS numbers from my previous elog are above for quick reference.

Case 1: Keep existing G&H mirror, flip it back the right way, and order new PR3/SR3. 

  • Spec PR3 to be concave with RoC 600 +/- 50m
  • This means the TMS in the PRC is in the range 1.4 MHz - 1.6 MHz [see this plot]
  • The mode matching efficiency for the PRC is > 98.5% [see this plot]
  • The TMS in the SRC is in the range 1.6 MHz - 1.8 MHz [see this plot]
  • Mode matching efficiency for SRC is > 98.5% [see this plot]
  • PRG between 34-38, depending on uncertainty in measurement of RoC of existing G&H mirror [see Attachment #1, added Nov 11 2016]

Case 2: Order two new sets of folding mirrors

  • Spec PR3/SR3 to be flat - for purposes of simulation, let's make it concave with RoC 10 +/- 5 km
  • Spec PR2/SR2 to be concave with RoC 1500 +/- 500m
  • The TMS in the PRC is between 1.7 MHz and 1.85 MHz [see this plot]
  • Mode matching efficiency is >98.5% in the PRC [see this plot]
  • TMS in the SRC is between 1.7 MHz and 2 MHz [see this plot]
  • Mode matching efficiency >99.0% in the SRC [see this plot]

At first glance, it looks like the tolerances are much larger for Case 2, but we also have to keep in mind that for such large RoCs in the km range, it may be impractical to specify as tight tolerances as in the 100s of metres range. So these are a set of numbers to keep in mind, that we can re-iterate once we hear back from vendors as to what they can do.

For consolidation purposes, here are the aLIGO requirements for the coatings on the RC folding mirrors: PR2, PR3, SR2, SR3

Attachment 1: PRG.pdf
PRG.pdf
  12416   Wed Aug 17 08:47:43 2016 ericqUpdateSUSAir-bake finished

I turned off the air bake oven at 8:45AM. I'll leave the optics alone for a bit while it cools.

  12415   Tue Aug 16 21:54:27 2016 gautamUpdateSUSAir-bake - IN PROGRESS

I put in both ETMX and ETMY into the air-bake oven at approximately 8.45pm tonight. They can be removed at 8.45am tomorrow morning. 


  • Given that we had previously melted a thermocouple in this oven, and there have been no high temperature bakes in it since, we ran the oven at 100C for about 3 hours in the afternoon
  • After that, I left the oven door open for an hour for the interior to return to room temperature
  • I then re-connected the controller (which doesn't seem very precise, it pulses the AC power to the oven in order to control the temperature), and dialled the oven back down to heating level 4, which is what Bob had it set at. I then waited for a couple of hours for the oven to reach ~34C
  • Before putting the optics in, I gave the inside of the oven a quick wipe with a clean wipe, and palced a layer of Al foil on the bottom of the oven
  • The optics are sitting on their donuts (see Attachment #1) - the copper wire elevates the optic+donut slightly and provides a path for air flow
  • ETMY was drag wiped with acetone+isopropanol prior to baking (to remove acetone stains from soaking to remove epoxy residue
  • We will of course be cleaning the optics with first contact prior to re-installation in the vacuum chambers
  • I am not sure what the extra cylindrical piece in there is, but Bob advised me to leave it in there so that's what I did
  • I've observed the temperature over ~2hours since I first put it in, and the oven/controller isn't going bonkers, so I'm trusting the controller and leaving for the night
Attachment 1: IMG_3005.JPG
IMG_3005.JPG
  12414   Tue Aug 16 16:38:00 2016 gautamUpdateCOCRC folding mirrors - Numerical review

Here are the results for case 2: (flat PR3/SR3, for purpose of simulation, I've used a concave mirror with RoC in the range 5-15km, and concave PR2/SR2 - I've looked at the RoC range 300m-4km).

  • This is where we order two new sets of mirrors, one for use as PR2/SR2, and the other for use as PR3/SR3.
  • RoC of flat PR3/SR3 in simulation explored in the range 5km-15km (concave)
  • RoC of concave PR2/SR2 in simulation explored in the range 300m-4km (concave)

Attachment #1: Mode matching between PRC cavities and arm cavities with some contour plots

Attachment #2: Mode matching between SRC cavities and arm cavities with some contour plots

Attachment #3: Gouy phase and TMS for the PRC. I've plotted two sets of curves, one for a PR3 with RoC 5km, and the other for a PR3 with RoC 15km

Attachment #4: Gouy phase and TMS for the SRC. Two sets of curves plotted, as above.


Hopefully EricG will have some information with regards to what is practical to spec at tomorrow's meeting.


EDIT: Added 9pm, 16 Aug 2016

A useful number to have is the designed one-way Gouy phase and TMS for the various cavities. To calculate these, I assume flat folding mirrors, and that the PRM has an RoC of 115.5m, SRM has an RoC of 148m (numbers taken from the wiki). The results may be summarized as:

Cavity One-way Gouy phase [rad]           TMS [MHz]           
PRX 0.244 1.730
PRY 0.243 1.716
SRX 0.197 1.743
SRY 0.194 1.717

So, there are regions in parameter space for both options (i.e. keep current G&H mirrors, or order two new sets of folding mirrors) that get us close to the design numbers...

Attachment 1: PRC_consolidated.pdf
PRC_consolidated.pdf
Attachment 2: SRC_consolidated.pdf
SRC_consolidated.pdf
Attachment 3: GouyPRC.pdf
GouyPRC.pdf
Attachment 4: GouySRC.pdf
GouySRC.pdf
  12413   Tue Aug 16 11:51:43 2016 gautamUpdateCOCRC folding mirrors - Numerical review

Summary of roundtable meeting yesterday between EricG, EricQ, Koji and Gautam:

We identified two possible courses of action.

  1. Flip the G&H mirror (PR2/SR2) back such that the (convex) HR face is the right way round. We want to investigate what are the requirements on a new PR3/SR3 optic that will guarantee cavity stability and also give good mode matching.
  2. Order two new sets of mirrors (i.e. replace all 4 folding mirrors). In this case, we want to spec a flat (how flat is reasonable to specify? EricG will update us) PR3/SR3, and design a PR2/SR2 with some concavity that will guarantee cavity stability in the event PR3/SR3 deviates from flatness (but still within what we spec). The choice to make PR3 as close to flat as possible is because the angle of incidence in our arrangement means that any curvature on PR3 dominates astigmatism.

I have done some calculations to evaluate the first alternative. 

  • Based on yesterday's preliminary discussion, we felt it is not reasonable to spec mirrors with RoC > 4km (sag of ~80nm). So I restrict my analyses to the range 300m-4km
  • Koji has a measurement of the phase maps for the G&H mirrors. The measured curvature is ~-500m. In my simulations, I've tried to allow for error in this measurement, so I look at the range -450m to -700m for the G&H mirror.
  • The Gouy phase analysis suggests we should look for an RoC of +500m (concave) for the new PR3/SR3 to have a TMS of ~1.5 MHz. Anything flatter (but still concave) means the TMS gets smaller.
  • The mode-matching in this region also looks pretty good, between 98% and 99%
  • I will post results of the analysis for the second alternative here for comparison

Something else that came up in yesterdays meeting was if we should go in for 1" optics rather than 2", seeing as the beam spot is only ~3mm on these. It is not clear what (if any) advantages this will offer us (indeed, for the same RoC, the sag is smaller for a 1" optic than a 2").


Attachments:

Attachment #1: Mode-matching maps between PRX and Xarm cavities, PRY and Yarm cavities with some contours overlaid.  

Attachment #2: Mode-matching maps between SRX and Xarm cavities, SRY and Yarm cavities with some contours overlaid. 

Attachment #3: Gouy phase calculations for the PRC

Attachment #3: Gouy phase calculations for the SRC

 

Attachment 1: PRC_consolidated.pdf
PRC_consolidated.pdf
Attachment 2: SRC_consolidated.pdf
SRC_consolidated.pdf
Attachment 3: GouyPRC.pdf
GouyPRC.pdf
Attachment 4: GouySRC.pdf
GouySRC.pdf
  12412   Tue Aug 16 08:04:49 2016 SteveUpdateSUSETMY had Aluminum bridge

The SOS ETMY tower had and Aluminum bridge. How is it possible that this was true? Is SS better than Al for some quantitative scientifc reason ???

Their weight ratio as measured  =  922 / 307 g = 3

Destinies:  SS 304 / Al 6061  >  0.289 / 0.098  [ lb /in3 ] =  2.94

Attachment 1: AlBridge.jpg
AlBridge.jpg
  12411   Mon Aug 15 18:28:15 2016 gautamUpdateSUSAir-bake preparation

I assume that we are prepared to live with the pitch bias situation of ETMY (i.e. we can achieve a configuration in which there is some pitch bias to the coils, and the OSEMs are inserted such that the PD outputs are half their maximum value). Or at least that we don't want to go through the whole standoff-regluing procedure for ETMY as well.

So today I took the optic out, and began to make some preparations for the air bake.

  • Both optics are now sitting in their respective metal donuts. 
  • How do we want to bake the optics? Bob has said he has prepared the oven for this bake, and that he has configured the temperature controller to a setpoint of 34C, and a ramp time of 2 hours to reach that temperature from lab temperature (we should check this before putting the optics in there with our independent temperature sensor - also, he is away for the week now so we can't get his input on any of these). But what about the actual logistics of how the optics are going to be housed? Specifically:
    • Do we want the donut to sit on some sort of tray? Presumably it is not ideal to have the HR surface in close proximity to the oven floor? 
    • Does the oven need any special cleaning?
    • Do we cover the donut+optic setup with a glass jar? If we do, any particles we eject off the optic can't escape the confines of the bowl, and if we don't, detritus from elsewhere may settle on the optic?
    • How long do we want this bake to last? 24hours? 48 hours? Bob didn't have an answer when I asked him earlier in the afternoon...
  • I also removed the suspension block from the top of the towers of both ETMX and ETMY, so that Steve could work on sanding them before we acetone-wipe and bake the towers themselves.
    • It was very apparent that the weights of the two pieces were largely different (ETMY suspension block ~350g, ETMX suspension block ~960g), even though they have the same physical dimensions.
    • Investigation into why this was yielded nothing conclusive. But Steve and I think that the ETMY suspension block is made out of Aluminum rather than SS, which would explain why the wire grooves seem deeper in the ETMY piece than the ETMX piece. It is worth noting that the specification calls for SS and not aluminum. But the top piece of the ETMY suspension (and indeed the old ETMX suspension) looks different from the specification, in that they don't have tapped holes for the secondary wire clamps (see Attachment #1).
    • I'm not sure if this is important, but it is worth noting. Steve and I also checked the remaining suspension towers. We think that ITMY, BS, SRM and PRM have the correct (to specification) suspension block. We couldn't get a look at ITMX and didn't want to take the door off. So ETMY (and possibly ITMX) will be the only suspension(s?) with a different suspension block.
  • Steve's sanding efforts did not go ideally.
    • He was successful in removing the wire grooves.
    • But the sharp edge which is supposed to clamp the wire seems to have been rounded a little bit (see Attachment #1). 
    • Overall, the section that we was sanded looks lower (i.e. its like we've dug a small channel into the plane of the suspension block)
    • Given that we suspect the ETMY suspension block is Aluminum, it is likely that attempting to sand it will yield an even deeper channel.
  • Do we want to bake the suspension towers in the large baking oven? Presumably we don't want to bake the optics with anything else. But does the large oven need any special cleaning before we stick the towers in there?
  • ETMY has some acetone marks on it. I will try and have this removed by drag wiping with more acetone and isopropanol prior to the bake tomorrow. Anyways we will first-contact clean the HR (and AR) sides after the bake before installing the optic.

In summary, the questions that remain (to me) are:

  1. Are we okay using an Al suspension block?
  2. How perfectly do we want wire grooves from prior suspensions removed? It looks like sanding doesn't work well, do we want to consider sending this into the shop?
  3. Baking logistics, as described above.

I think we can start the baking of the optics tomorrow. The timeline for the suspension towers is unclear, depends on how we want to deal with the sanding dilemma.

Attachment 1: IMG_6816.JPG
IMG_6816.JPG
  12410   Mon Aug 15 14:34:33 2016 ericqUpdateSUSSOS sus wire ordered

We have indeed seen numerous tarnished/rusty points along the wires, and just tried to choose lengths free of any of these. I wonder if this can explain the brittleness/ease with which we've been breaking it. My feeling is that we should use the newer wire if feasible.

  12409   Mon Aug 15 14:29:32 2016 SteveUpdateSUSSOS sus wire ordered

The wire will arrive in 1-2 weeks. It is a new production. Brad Snook of Ca Fine Wire was suprised that we are still using the 13 years old wire.  Oxidation is an issue with iron contained steel wire.

He would not give me a shelf life time on it. He recommended to check the strenght of it before usage. It passed with safety factor of 2 just recently.

In the future we'll store the new spool in oxigen free nitrogen environment..

Quote:

0.0017" OD., 500ft steel music wire ordered. Pictures of the existing roll are below. It will on 9" OD. spool too.

 

 

  12408   Mon Aug 15 12:23:56 2016 PrafulUpdatePEMMic Self Noise

I didn't have a separate training set and data set, so I think that's why the graphs came out looking too good. The units on the graphs are also incorrect, I was interpreting PSD as ASD. I haven't been able to get my Wiener filtering code working well- I get unreasonable subtractions like the noise being larger than the unfiltered signal, so Eric showed me this frequency-dependent calculation described here: https://dcc.ligo.org/LIGO-P990002

This seems to be working well so far:

freq1.pdf

freq2.pdf

freq3.pdf

Here's all the plots on one figure:

frequency_dependent.pdf

Let me know if this looks believable.

Quote:

Seems to good to be true. Maybe you're over fitting? Please put all the traces on one plot and let us know how you do the parameter setting. You should use half the data for training the filter and the second half for doing the subtraction.

 

Attachment 1: freq1.png
freq1.png
Attachment 2: freq1.pdf
freq1.pdf
Attachment 4: freq2.pdf
freq2.pdf
Attachment 6: freq3.pdf
freq3.pdf
Attachment 8: frequency_dependent.pdf
frequency_dependent.pdf
  12407   Sat Aug 13 18:25:22 2016 gautamUpdateCOCRC folding mirrors - Numerical review

This elog is meant to summarize my numerical simulations for looking into the effects of curvature on the RC mirrors. I've tried to go through my reasoning (which may or may not be correct) and once this gets a bit more refined, I will put all of this into a technical note.


Motivation: 

  • Both the G&H (PR2, SR2) and Laseroptik (PR3 SR3) are convex on the HR side with RoCs of approximately -600m and -700m (though as stated in the linked elog, I'm not actually sure if there are measurements of this number) EDIT AUG15: There are measurements for the Laseroptik mirrors here
    GV April 8 2017: This elog by Jenne suggests that the installed PR2 has an RoC of approximately -700m. Koji has uploaded the phase map data for the RC TT mirrors to
    /users/public_html/40m_phasemap/40m_TT and 
    /users/public_html/40m_phasemap/40m_TT2. The G&H mirror data seems to be in the former folder, and it looks like there are two mirrors, one with RoC of ~ -700m and the other with RoC of ~ -500m. Does this mean PR2 has RoC -700m and SR2 has RoC -500m?
  • As a result, both the PRC and SRC were close to instability
  • By flipping the folding mirrors, the instability has been mitigated, but at the expense of the non-ideal situation where the AR coated side and the substrate are now inside the recycling cavity
  • We would like to order some new folding mirrors. In order to avoid receiving convex mirrors from the vendor, we want to specify a concave curvature for the HR side
  • The aim of this investigation is to look at how concave we should make these mirrors, because although the cavity stability improves with concavity of the HR side, possible disadvantages of having too convex mirrors are:
    • ​Mode-mismatch between the recycling cavities and the arms
    • Astigmatism

The study:

  • I've built a Finesse model for the 40m, which has been used for all the numerical studies quoted here
  • In constructing this Finesse model, I've used the following sources to specify various paramaters:
    • ​RoCs, R, T and physical dimensions of 4 test-masses, PRM, SRM and BS: Core optics wiki page
    • Losses - arm losses from Yutaro's measurements in elog11857 and elog11818 (distributed equally between ITM and ETM). For other optics, a generic value of 25ppm was used
    • "Ideal" lengths for our current modulation frequency were used for the various cavities (37.795m for the arms, 6.753m for PRC, 5.399 for SRC)
    • The folding mirrors (PR2, PR3, SR2, SR3) are initialized as flat in the model
  • I performed some low-level checks (e.g. arm linewidth, PRC FSR etc) to check that the model was sensible
  • I then proceeded to investigate the effects of curvature on the folding mirrors. Specifically, I investigated the following:
    • What is the mode mismatch between the recycling cavity mode and the arm as a function of the RoC of the folding mirror?
    • What is the effect of the RoC of the folding mirrors on the round-trip gouy phase accumulated (and hence the transverse mode spacing) in the recycling cavities?
  • For now, the parameter space explored is from 300m concave to 1000m concave. An RoC of 1km for a 2" optic corresponds to a sag of ~0.3 microns. I will explore the 1km-10km concave space and update the results shortly

Results:

  • Attachments #1 and #2 show the mode mismatch between the recycling cavity and the arm for various curvatures. The colorbars have been normalized to span the same range in all the plots
  • For both the PRC and the SRC, if we have folding mirrors with an RoC of 1000m concave, we will have a mode mismatch of 2-3%. The number gets worse the more convex the mirror
  • Attachments #3 and #4 show the one-way accumulated Gouy phase. Here, I have varied the curvature of the folding mirrors along a specific axis at a time (i.e. I've assumed that the folding mirrors are identical). I've also added the transverse mode spacing as a second y-axis. I have yet to check how these numbers compare with the linewidth of the 00-mode for the various fields, but for 1km concave folding mirrors, the TMS is in the region of 2MHz 

To do:

  • I will extend the range of RoCs explored to 10km concave and post results - but I will have to check with EricG to make sure that it is feasible for us to specify curvatures in this range
  • I was trying to use the RT gouy phase as calcluated by my Finesse simulations to plug into some analytical expressions to try and generate plots like this for various RoCs of the folding mirrors, but if the TMS calculations suffice, I will abandon these efforts
  • What are the other specifications we need to worry about before placing an order? Some thoughts from Rana's earlier elog:
    • The coatings need to be dichroic to allow extraction of the green beam (but only PR3/SR3 is currently dichroic?)
    • Wedge angle on the AR side?
  • Are there any other obvious sanity checks I should carry out?

 

Attachment 1: PRX_consolidated.pdf
PRX_consolidated.pdf
Attachment 2: SRX_consolidated.pdf
SRX_consolidated.pdf
Attachment 3: Gouy_PRC.pdf
Gouy_PRC.pdf
Attachment 4: Gouy_SRC.pdf
Gouy_SRC.pdf
  12406   Fri Aug 12 21:26:28 2016 ranaUpdatePEMMic Self Noise

Seems to good to be true. Maybe you're over fitting? Please put all the traces on one plot and let us know how you do the parameter setting. You should use half the data for training the filter and the second half for doing the subtraction.

  12405   Fri Aug 12 19:13:25 2016 PrafulUpdateElectronicsMic Self Noise

I used the Wiener filtering method described by Ignacio and Jessica (https://dcc.ligo.org/DocDB/0119/T1500195/002/SURF_Final.pdf and https://dcc.ligo.org/public/0119/T1500194/001/Final_Report.pdf) and got the following results:

mic1_wiener.pdf

mic2_wiener.pdf

mic3_wiener.pdf

The channel readout has a gain of 0.0005 and the ADC is 16-bit and operates are 20V. The channel also reads the data out in Pa. I therefore had to multiply the timeseries by 1/0.0005=2000 to get it in units of counts and then by (20 Volts)/(2^16 counts) to get back to the original signal in volts. The PSDs were generated after doing this calibration. I also squared, integrated, and square rooted the PSDs to get an RMS voltage for each microphone as a sanity check:

Mic 1: 0.00036 V

Mic 2: 0.00023 V

Mic 3: 0.00028 V

These values seem reasonable given that the timeseries look like this:

timeseries_elog.pdf

 

 

Attachment 4: mic1_wiener.pdf
mic1_wiener.pdf
Attachment 5: mic2_wiener.pdf
mic2_wiener.pdf
Attachment 6: mic3_wiener.pdf
mic3_wiener.pdf
Attachment 7: timeseries_elog.pdf
timeseries_elog.pdf
  12404   Fri Aug 12 14:37:34 2016 SteveUpdateSEIworking seismometers as they are

2.1 mag earth quake in Norhten Ca

Our seimometers need professorial centering. Related electronics must be checked too.

Quote:

The saga has started here  We have to give credit to the Boss who fixed it. The seismometers themself are not labeled yet.

Atm6  added on 8-12-2016   EX needed to be centered

Thanks to Max for the nice plost at summery pages

 

 

Attachment 1: eq_2.1m_Geysers_CA.png
eq_2.1m_Geysers_CA.png
Attachment 2: eq2.1m_GeysersCA.png
eq2.1m_GeysersCA.png
Attachment 3: oscillations.png
oscillations.png
  12402   Thu Aug 11 17:30:05 2016 PrafulUpdateElectronicsMic Amplifier

The results of my first huddle test were not so good- one of the signals did not match the other two very well- so I changed the setup so that the mics would be better oriented to receive the same signal. Pictures of the new setup are attached.

I also noticed some problems with one of my microphones so I soldered a new mic to bnc and switched it out. Just judging from Dataviewer, the signals seem to be more similar now. I'll be taking data for another few hours to confirm.

  12401   Thu Aug 11 11:56:40 2016 gautamUpdateSUSETMY re-suspended
Quote:

How much pitch bias do you need in order to correct this pitch misalignment?
That may give you the idea how bad this misalignment is.

I needed to move the pitch slider on the IFO align screen to -2.10 (V?) from 0 to get the HeNe spot to the center of the iris. The slider runs from -10V to 10V, so this is something like 10% of its range. I am not sure if it means anything, but the last saved backup value of this pitch slider was -3.70. Of course, application of the bias will affect all the coils, and when the optic is pitch balanced, the lower magnets are a little too far out and the upper magnets are a little too far in (see Attachment #1), as we expect for a downward pitch misalignment to be corrected. I suppose we can iteratively play with the coil positions and the bias such that the coils are centered and we are well balanced (maybe this explains the old value of -3.70). 

I also checked that the side magnet can completely occlude its PD. With the damping on, by pushing the coil all the way in, the output of the side PD went down to 0.

Attachment 1: pitchBalancingWithBias.PDF
pitchBalancingWithBias.PDF
  12400   Thu Aug 11 11:51:38 2016 PrafulUpdateComputer Scripts / ProgramsSummary Pages

The summary pages have been updated with the new naming seismometer channel naming conventions. Here's a link to them working on my own page: https://ldas-jobs.ligo.caltech.edu/~praful.vasireddy/1154908817-1154909717/pem/seismic/
 

Let me know if the actual pages aren't working when they come back online or if there's something that needs to be changed.

  12399   Thu Aug 11 11:09:52 2016 Max IsiUpdateGeneralSummary pages status
This problem has been fixed.

> Summary pages are currently empty due to a problem with the code responsible for locating frame files in the cluster. This should be fixed soon and the
> pages should go back to normal automatically at that point. See Dan Kozak's email below for details.
>
>
> Date: Wed, 10 Aug 2016 13:28:50 -0700
> From: Dan Kozak <dkozak@ligo.caltech.edu>
>
>
> > Dan, maybe it's a gw_data_find problem?
>
> Almost certainly that's the problem. The diskcache program that finds
> new data died on Saturday and no one noticed. I couldn't restart it,
> but fortunately it's author just returned from several weeks vacation
> today. Smile He's working on it and I'll let you know when it's back up.
>
> --
> Dan Kozak
> dkozak@ligo.caltech.edu
  12398   Thu Aug 11 00:20:41 2016 KojiUpdateSUSETMY re-suspended

How much pitch bias do you need in order to correct this pitch misalignment?
That may give you the idea how bad this misalignment is.

  12397   Wed Aug 10 23:45:03 2016 gautam UpdateSUSETMY re-suspended

Summary:

  • ETMY has been re-suspended
  • Reglued magnets (and also those that weren't knocked off) quite well with OSEM coils (see attachments)
  • Pitch balance is off by ~2.8mrad (8mm over 1.5m lever arm) after inserting and centering OSEMs
  • The same damping scheme used during the ETMX re-suspension process works reasonably well with ETMY as well

Details:

  • I suspected that I had not tightened the wire clamp enough yesterday, and that the wire had slipped once the winches were removed
  • Steve and I looked into the torque wrench situation today, and I realised that I had not been using the torque wrench correctly. What I thought were clicks indicating that the set torque has been reached was in fact just the sound the piece makes when going the opposite way relative to the direction set by the clip on the torque wrench. Anyways, the point is that while I thought I was tightening the screws with ~1.3Nm of torque, what was actually being applied was much less (although I don't have a good way to quantify how much less)
  • So today I put the winches back on top of the tower, and winched the optic back up to the correct height using the ususal scribe line + microscope prescription
  • I then tightened the wire clamp by hand. This is obviously not very repeatable, but it will have to do until we get a torque wrench with the correct range
  • This seems to have done the trick - I did the tightening shortly after lunch, and after ~10 hours, there is no evidence of any wire sag
  • I then proceeded to insert the OSEMs, first not all the way in to check the clearance available to the magnet, and once I was satisfied there was no danger of knocking anything off, went ahead and inserted the coils till the PD readouts were approximately half of the maximum (i.e. fully un-occluded) values. I used the OSEM coils originally on the ETMY tower, but all the other readout and drive electronics in the signal chain (satellite box included) belong to the ETMX setup (so as to avoid any cable routing over 80m from the Y end to the cleanroom). After some adjustment of the OSEM holding plates, I was able to center the magnets relative to the coils 
  • The tower only allows for a side OSEM to be inserted on one side. The other side does not have a threaded hole for a set screw. So we are forced to use the reglued magnet and not the side magnet that was not knocked off. By eye, it looks like the magnet may never completely occlude the LED, but the Striptool trace I was using to monitor the output of the PD did not yield any conclusive evidence. The optic was moving around a lot and I did not perform this check after turning the damping on
  • I was able to damp the optic as well as we were able to damp ETMX on the clean bench (with the HEPA turned OFF). I had to turn the YAW gain down from 100-->75 to avoid some oscillations 
  • I then proceeded to check the pitch balance with the HeNe. The spot is low on an iris 1.43m away by ~8mm, which corresponds to a pitch misalignment of ~2.8mrad. I am not sure what to make of this - but perhaps its not unreasonable that we see this? Is there any record of what fine pitch balancing was achieved when the optic was put together back in 2010? This is also very sensitive to how far in/out the OSEM coils are, and though I've tried to center the coils as best as I can, I obviously have not done a perfect job...

What's next?

  • Is the observed pitch imbalance a deal breaker? If so, I guess we need to re-glue a standoff? 
  • Are we willing to accept the side OSEM situation? (Tomorrow, I need to do a check to see what, if any, dynamic range we lose, with the damping on)
  • If both the above are not problems we need to worry about, then:
    • ETMY + ETMX -------> Vacuum bake on 22nd August (? - Bob also told me earlier today that he will try and put in some old turbo pump next week, and if that works, we could possibly get in the queue even before the 22nd)
    • ETMY tower -------> Steve for sanding and removing wire grooves -------> Air bake
    • ETMX tower -------> Air bake (provided the latest round of wire tightening has not left any grooves in the top piece of the tower, if it has, this needs to be cleaned up too)
    • Some lengths of SOS wire (for re-suspending optics after bake) -------> Air bake

Attachments:

Attachment #1: Striptool trace showing all OSEM coils have been pushed in till the PD readout is approximately half the fully open value

Attachment #2: Pitch balance is off by ~2.8mrad (the Iris center is 5.5" above the table)

Attachment #3: UR magnet

Attachment #4: UL magnet

Attachment #5: LR magnet

Attachment #6: LR magnet

Attachment #7: SD magnet

Attachment 1: ETMY_OSEMStrip.PDF
ETMY_OSEMStrip.PDF
Attachment 2: IMG_2998.JPG
IMG_2998.JPG
Attachment 3: IMG_3000.JPG
IMG_3000.JPG
Attachment 4: IMG_3001.JPG
IMG_3001.JPG
Attachment 5: IMG_3002.JPG
IMG_3002.JPG
Attachment 6: IMG_3003.JPG
IMG_3003.JPG
Attachment 7: IMG_3004.JPG
IMG_3004.JPG
  12396   Wed Aug 10 19:37:08 2016 gautamUpdateElectronicsMic Amplifier

In order to help Praful do his huddle test, I have temporarily arranged for the outputs of the 3 channels he wants to monitor to be acquired as DQ channels at 2048 Hz by editing the C1PEM model. No prior DQ channels were set up for the microphones. Data collected overnight should be sufficient for Praful's analysis, so we can remove these DQ channels from C1PEM before committing the updated model to the svn. There is in fact a filter that is enabled for these microphone channels that claims to convert the amplified microphone output to Pascals, but it is just a gain of 0.0005. 

In the long term, once we install microphones around the IFO, we can update C1PEM to reflect the naming conventions for the microphones as is appropriate.

  12395   Wed Aug 10 18:10:26 2016 PrafulUpdateElectronicsMic Amplifier

I set up 3 of my circuits in the interferometer near MC2 to do a huddle test. I have the signals from my microphones going into C1:PEM-MIC_1_IN1, C1:PEM-MIC_2_IN1, and C1:PEM-MIC_3_IN1. These are channels C17-C19. Here are some pictures of my setup:


I'll likely be collecting data from this for a couple of hours. Please don't touch it for now- it should be gone soon. There are some wires running along the floor near MC2 as well.

  12394   Wed Aug 10 17:30:26 2016 Max IsiUpdateGeneralSummary pages status
Summary pages are currently empty due to a problem with the code responsible for locating frame files in the cluster. This should be fixed soon and the
pages should go back to normal automatically at that point. See Dan Kozak's email below for details.


Date: Wed, 10 Aug 2016 13:28:50 -0700
From: Dan Kozak <dkozak@ligo.caltech.edu>


> Dan, maybe it's a gw_data_find problem?

Almost certainly that's the problem. The diskcache program that finds
new data died on Saturday and no one noticed. I couldn't restart it,
but fortunately it's author just returned from several weeks vacation
today. Smile He's working on it and I'll let you know when it's back up.

--
Dan Kozak
dkozak@ligo.caltech.edu
  12393   Wed Aug 10 17:20:16 2016 LydiaUpdateGeneralSeismometer channel names changed

[ericq, Lydia]

We changed the seismometer channel names from, e.g. C1:PEM-SEIS_GUR1_X to C1:PEM-SEIS_EY_X.

  • GUR1 referred to the seismometer at the Y end, GUR2 referred to the seismometer at the X end, and STS1 referred to the seismometer at the vertex. These have been renamed EY, EX, and BS respectively in all channel and filter names. 
  • The models for c1pem and c1oaf were changed in Simulink. The DAQ boxes were also updated with the new names. The script which forcibly renames channels saved to disk was edited to no longer refer to the GUR1, GUR2 etc channels. Going along with what Rana suggested we decided not to change the names of the renamed channels this way when saving, so the data saved from the seismometers can be found under e.g. C1:PEM-SEIS_EY_X_OUT_DQ. 
  • The filter files generated by Foton were changed to reflect the new channel names. 
  • We compiled the changes to the models and restarted the models on the relevant machines (c1lsc for the c1oaf model and c1sus for the c1pem model). c1sus_aux was down so we manually restarted it to turn off the watchdogs as a precaution before putting too much strain on c1sus. 
  • The MEDM screens now show the correct information, relabeled with the new names under PEM-RMS. 
  • The striptool display projected on the wall now shows the appropriate C1:PEM-RMS_BS channels and has been renamed to "SeismicRainbowBS.strip" 
  • We verivied that the new channels can be accessed live and the data from the DQ channels is saved to disk. 
  • After the changes were complete, we attempted to commiting to svn (the commit also included bringing the MEDM screen files into version control.) However the svn server was taking a long time to respond, so we will try again tomorrow to commit the file changes. 
  • There are still some lefotver unused channels with name sinculding STS_2 and STS_3 that refer to seismomters we no longer use. We left these alone. 

Summary of new channel names:

C1:PEM-RMS_{BS, EX, or EY}_{X, Y, or Z} followed by the same filtering options as before, e.g. _BP_1_3_OUT

C1:PEM-SEIS_{BS, EX, or EY}_{X, Y, or Z}_{EXC, IN1, IN2, OUT, or OUT_DQ} 

C1:OAF-WIT_{BS, EX, or EY}_{X, Y, or Z}_{EXC, IN1, IN2, or OUT} 

 

  12392   Wed Aug 10 15:34:24 2016 SteveUpdateSUS6 in-lbs torque driver for wire clamp screw

The 7.5 in-lb of Wiha seems at the upper end of torque range for a 4-40 SS screw

Wiha 28502 ordered with range 5 -10 in-lb for silver plated 4-40 screws

Do not trust the Venzo torque wrench under 2 Nm ! It miss lead me.

Recommended torque values for silver-plated fasteners are here. For aLIGO we use the guidelines in T1100066-v6, This doc is posted in 40m wiki under Mechanics also.

So, we'll use 6 in-lbs  on silver plated 18-8 stainless steel socket head cap screw 4-40 x 3/8 into SS tower bridge.

Please replace these clamp screws every time if they were tightened without a torque wrench.

Quote:

New Wiha 28504 torque wrench for SOS wire clamping. It's range 7.5  - 20 in-lb in 0.5 steps [ 0.9 - 2.2 Nm ] Audible and perceptible click when the pre-set torque has been attained at ±6% accuracy.  

The new ETMX sus wire torqued to ~ 11.5 in-lb [1.3 Nm ]

Quote:

Gautam and Steve,

The clamp's left side was jammed onto the left guide pin. It was installed slit facing left. Gautam had to use force to remove it. The clamp should move freely seating on the guide rods till torque aplied. Do not move on with the hanging of optic with a jammed clamp. Fix it.

Never use force as you are hanging - aligning optic. The clamp is in the shop for resurfacing and slit opening.

 

 

 

  12391   Wed Aug 10 10:32:44 2016 SteveUpdateSEIworking Guralps as they are

The saga has started here  We have to give credit to the Boss who fixed it. The seismometers themself are not labeled yet.

Atm6  added on 8-12-2016   EX needed to be centered

Thanks to Max for the nice plost at summery pages

 

Attachment 1: InterfboxOutpFixed.png
InterfboxOutpFixed.png
Attachment 2: GuralpsCoherences.png
GuralpsCoherences.png
Attachment 3: GurInterfBox.jpg
GurInterfBox.jpg
Attachment 4: EX.jpg
EX.jpg
Attachment 5: EY.jpg
EY.jpg
Attachment 6: EX_EY.png
EX_EY.png
  12390   Wed Aug 10 03:08:03 2016 gautam UpdateSUSETMY patch-up

[lydia, gautam]

Rana felt it was alright to use the wire clamp and suspension cage in its existing condition for checking the ETMY magnet-OSEM coil alignment. So we set about trying to re-suspend ETMY. The summary of our attempts:

  • Transferred optic from magnet gluing rig to the suspension cage
  • Adjusted bottom EQ stops till the scribe lines on both sides were at 5.5" as verified with the microscope
  • Looped cleaned length of wire around optic, attached free ends to winches, placed the wires under light tension by finger-pulling the slack out
  • Lowered the bottom EQ stops
  • Winched the optic to the right height
  • Clamped the wire with the only wire clamp on this variant of the suspension cage. We used the same torque wrench at the same torque setting as was successful for ETMX. But after removing the winches, and releasing the face EQ stops, the optic seems to have sagged a lot - it now touches all the bottom EQ stops, and the more I lower it, the more it seems to come down. Perhaps it is the effect of the wire grooves in the cage, or that the wire-clamp itself is slightly different from the piece used on the ETMX cage, but 1.3Nm of torque doesn't seem to have tightened the wire clamp sufficiently
  • We can still probably salvage the situation by re-attaching the winches to the top of the cage, setting the optic to the right height again, and clamping the wire clamp with more torque (as this is just a check to see that the reglued magnet configuration is compatible with the OSEM coil positions on the cage). Before air baking the cage, we will have the old wire grooves removed, and then suspend the optic with a fresh loop of wire after the bake
  • We could not check the magnet-OSEM alignment because of the slipping of the wire through the clamp. We decided against pushing on tonight
  • Optic is currently in the cage, resting on the bottom EQ stops and with all face EQ stops within 1mm of the optic. The OSEM coils have not been inserted into the holders

Regarding the vacuum bake of the optics: why do we want to do this again? Koji mentioned that the EP30-2 curing process does not require a bake, and there is also no mention of requiring a vacuum bake in the EP30-2 gluing guide. Is there any other reason for us to vacuum bake the optic?

  12389   Tue Aug 9 19:35:49 2016 ranaUpdateSEIboth Guralp seismometers are functioning and being acquired

After some cable swapping, we now have both Guralp seismometers running and the times series and spectra look similar to each other and motley blushhealthy.


Bean and I took a look at the whole situation today. Ben had nicely fixed the Dsub end of the EX cable (the EY one is still just a sad joke), After installing this newly fixed cable, we still saw no signals. There was some confusion in the control room about using the MED displays to diagnose seismometers: flickering MEDM values cannot be used for this. It would be like checking a pizza box temperature to determine if the pizza is any good.

  1. Although the +/- 12V LEDs on the front panel are dim, we confirmed that the produce 11.94V even when loaded with a seismometer. So its a LED circuit problem not a power problem.
  2. We were able to inject signals into the front panel with a breakout board and see them in DV for Input 1, but not Input 2.
  3. After Ben left, I kept poking around and found that the Guralp chassis output gets broken out into 3x3 BNC cables before going to the PEM BNC panel (and then on to the PEM ADC). This is where the problem was.
  4. The Input #3 BNC cables were connected to the long cables going to the 'GUR2' channels of the PEM. The Input#2 BNC cables were connected to some short BNC cables that were just hanging from the rack. So, somewhere during the debugging of the past N months, someone plugged this in wrong and didn't notice or forgot to switch it back. So all of the tests using DV or DTT or MEDM since that time have been invalid.

 

Tomorrow, Lydia is going to change all of the labels and channel names. The new names will be EX & EY to prevent this kind of huge waste of time with channel name swapping. That means no more illegal names with the label maker, Steve.

From the spectrum you can see that the EX seismometer (GUR2) is still not centered or at least its oscillating at 245 Hz for some reason. This should go away after some power cycling or recentering using the magic wand.

I noticed some anomalies in the mechanical setups at the ends:

  1. Some junk has been stored on top of the EX seismometer. Please never, even temporarily, store your power supplies, tools, or donutsfrown on top of the vibration sensitive sensors. Just put it on the floor and improve your carma.
  2. The EY seismometer has some fishy wires being fished between the can and the rubber seal. This is verboten. That seal must be flush to prevent pressure fluctuations and wires in there will ruin the smooth contact permanently. Temperature sensor wires must go through the grantie block feed-through or else its pointless.
  3. The flimsy insulation on the EY seismo is waay toooo mickey mouse. Real thermal insulation should be done using the yellow foam that Jenne used for the seismo huddle test. This flimsy silvery stuff is OK for making hats and mittens and beer cozy's, but its not research grade foam.
Attachment 1: 1goodday.png
1goodday.png
  12388   Tue Aug 9 16:19:27 2016 SteveUpdateSUStorque driver for wire clamp

New Wiha 28504 torque wrench for SOS wire clamping. It's range 7.5  - 20 in-lb in 0.5 steps [ 0.9 - 2.2 Nm ] Audible and perceptible click when the pre-set torque has been attained at ±6% accuracy.  

The new ETMX sus wire torqued to ~ 11.5 in-lb [1.3 Nm ]

Quote:

Gautam and Steve,

The clamp's left side was jammed onto the left guide pin. It was installed slit facing left. Gautam had to use force to remove it. The clamp should move freely seating on the guide rods till torque aplied. Do not move on with the hanging of optic with a jammed clamp. Fix it.

Never use force as you are hanging - aligning optic. The clamp is in the shop for resurfacing and slit opening.

 

 

  12387   Tue Aug 9 15:50:30 2016 PrafulUpdateGeneralGuralp Cable

The Guralp cable has been reconnected and powered after having the connector changed out.
 

  12386   Tue Aug 9 15:27:57 2016 gautam UpdateSUSETMY patch-up

The pickle pickers came off nicely and both magnets seem to be glued on okay. The alignment of the face magnets look pretty good, but we will only really know once we suspend the mirror, check the pitch balance, and put in the OSEM coils.

I brought the ETMY suspension tower + OSEM coils out of the vacuum chamber into the cleanroom. Given that the old wire had a pretty sharp kink in it, I removed it with the intention of suspending the optic with a new length of wire. I noticed a few potential problems:

Attachment #1 - ETMY tower is different from ETMX tower: 

  • The ETMY suspension seems to be of an older generation - it does not have the the two secondary wire clamps. 
  • The top piece was attached to the body of the tower using non-silver-plated screws. Steve tells me this is the wrong type, and we can switch these out when we put it back together.
  • The wire clamp itself doesn't have much of a groove from the wire. But the wires have made asymmetric grooves in the tower itself (the left groove is deeper than the right as seen in Attachment #1), that are clearly visible. Should we get these grooves removed before attempting re-suspension? How do we want to remove it? Steve thinks the best option is to send it to the shop for milling, as there is hardly any room to rub sandpaper along the piece because of the pins, and these pins don't come out. 
  • Or do we just not care about these grooves for now, if we are planning to use new wire anyways after air-baking the towers? 
  • Steve thinks we should have a few spares of these top blocks handy (the latest version, with the secondary clamps), he wants to know if we should place an order for these (we already have 10 spare wire clamp pieces available for if/when we need them)

Attachment #2 - the base of the tower is significantly rusty:

  • A few wipes with an acetone soaked rag yielded quite a lot of rust
  • Steve thinks this is because the wrong type of stainless steel was used
  • Does this have to do with the cage being of an older variety? After a few vigorous wipes, no more rust came off, but the rusting process will presumably keep generating new rust? Is this a concern? Do we want to change this piece before putting the tower back in?

I am holding off on attempting to re-suspend the optic for now, until we decide if the old wire grooves need to be removed or not. If we are okay with re-using the same piece as is, or if we are okay with using sandpaper and not the machine shop to remove the grooves, I will resume the re-suspension process. 

Eric suggested another alternative, which is to use the old ETMX tower. I don't recall it being rusted, but this has to be checked again. The other problem of the wire-grooves would possibly still be an issue.


Regarding the vacuum bake of the ETMs, Bob tells us that the best case scenario we are looking at is September.

 

Attachment 1: IMG_2996.JPG
IMG_2996.JPG
Attachment 2: IMG_2997.JPG
IMG_2997.JPG
  12385   Tue Aug 9 13:53:57 2016 babbottUpdateSEIlong Guralp EX cable repaired on the D-sub side

I checked out the cable that I took from you, and all of the connections looked right.  The only thing I did notice was that some of the soldered wires on the 37-pin connector had gotten hot enough to melt their insulation, and potentially short together.  I cut off that connector, and left it on your desk to check out.  I put on a new connector, and checked the pinout.  If the Guralps still doesn't work, we'll have to check out other possibilities.

  12384   Tue Aug 9 00:44:43 2016 gautam UpdateSUSETMY patch-up

Summary:

Given that ETMX looks to be in good shape and the optic and suspension tower are ready for vacuum and air bakes respectively, I set about re-gluing the knocked off magnet of ETMY. In my previous elog, I had identified the knocked off magnet as the UL magnet. But in fact, it was the LR magnet that broke off. This is actually one of the magnets that was knocked off when Johannes was removing the optic from the vacuum chamber. I have edited the old elog accordingly. 

Step 1: Removing epoxy residue

  • I used the teflon+glass rig Steve put together for this purpose
  • After soaking for ~2 hours in acetone, I was able to remove approximately half of the ring residue by lightly pushing with a wipe.
  • The other half wouldn't budge so I let it soak for another 4 hours
  • After 6 hours of soaking, I was able to get all of the epoxy residue off - it doesn't simply dissolve in the acetone, I had to push a little with one of the cotton-tipped paddles in the cleanroom
  • I gave the portion exposed to acetone a quick drag wipe with isopropanol. I didn't spend too much time trying to clean the AR side given that we will be using first contact anyways.
  • I have not touched the HR side for now, even though a small portion of it was exposed to acetone. While cleaning the HR face with first contact, this portion can be inspected and cleaned if necessary

Step 2: Putting the optic in the magnet gluing jig

  • I transferred the optic to the magnet gluing jig
  • Given that we weren't touching any side magnets, I reasoned I did not have to go through the elaborate shimming routine to account for the wedge of the optic that we had to do in the recent past
  • However, I did not think to put a thicker teflon spacer on the lower side of the wedge, and as a result, I knocked off the UR magnet as well as the jig did not have sufficient clearance
  • Fortunately, the UR magnet came off cleanly, there was hardly any epoxy residue left on the optic. The UR magnet was NOT one of the magnets knocked off by Johannes while removing the optic from the vacuum chamber
  • I gave the area formerly occupied by the UL magnet 3-4 wipes with acetone and then 1-2 wipes with isopropanol
  • At this stage, I proceeded to re-insert the magnet-gluing jig. I used the two scribe lines on the outer side of the jig to fix the rotation of the jig, and used the remaining two attached face magnets to fix the overall position of the jig (by centering these magnets relative to the apertures on the jig). In order to center well, I had to unscrew the stuck silver plated screw on the jig by 1 turn
  • Having arranged the jig satisfactorily, I proceeded to remove epoxy residue off the dumbbell of the recently knocked off UL magnet using first a razor blade, then sandpaper and finally made some new grooves with a razor blade. I then cleaned the surface of the dumbbell to be in contact with the optic with isopropanol. All of this was done for the LR magnet two weeks ago right after it was knocked off

Step 3: Gluing the magnets

  • I prepared the magnets in the pickle pickers
  • I discarded 1 full squeeze of the epoxy after it reached the tip of the mixing fixture, and then extracted another full squeeze of the gun for mixing and gluing the magnets
  • I mixed the epoxy in an Al foil vessel for 3-4 minutes, and then placed a few drops on a piece of Al foil for a test bake at 200F for ~15 minutes
  • The test bake went well, so I proceeded to apply glue to the dumbbells and re-glue the magnets to the optic
  • The gluing was done around midnight, so we should be able to have a look at this post lunch tomorrow.

Provided the gluing goes well, the plan for tomorrow is:

  1. Bring ETMY suspension tower from the vacuum chamber to the cleanroom along with its OSEMs
  2. Suspend ETMY with a new length of wire (this should be much more straightforward than our ETMX exploits as both standoffs are already glued)
  3. Insert OSEMs, check that all 4 face magnets are well centered w.r.t. their coils and also that at least one side magnet is well aligned relative to its coil and can be used 
  4. If step 3 goes well, then ETMY is also ready for a vacuum bake. I guess we can also air bake the ETMY suspension tower, there's plenty of room in the oven
  12383   Mon Aug 8 10:35:12 2016 SteveUpdatePEMvent 78 day 38

 

Quote:

RGA background scan

Quote:

Vacuum Status: Chamber Open

All chamber annuloses are vented.  Vac Monitor screen is not communicating with gauges. The valve position indicator are working.

RGA is pumped by Maglev through VM2

 

 

 

Attachment 1: Vent38days.png
Vent38days.png
Attachment 2: vent78d40RgaBg.png
vent78d40RgaBg.png
ELOG V3.1.3-